In a survey, 205 people indicated they prefer cats, 160 indicated they prefer dots, and 40 indicated they don’t enjoy either pet. Find the probability that if a person is chosen at random, they prefer cats.

Explain in details.
Answer #1

In the survey,

People prefer cats = 205

People prefer dogs = 160

People prefer neither pets = 40

Number of sample in the survey = (205+160+40) = 405

The probability that if a person is chosen at random, they prefer cats

= (Number of people prefer cats)/(Total number of sample)

= 205/405 = 41/81